# Enhancing Agent Learning through World Dynamics Modeling

### v1

@@ 4.4 Qualitative Analysis
The correctness of the learned and verified dynamics is categorized as either correct or as mistakes due to confounders, in-domain hallucinations, or out-of-domain hallucinations. As illustrated in Table 3, in the example of defeating a zombie, the mistake attributed to confounders is primarily due to the simultaneous increase in health points at the step where the zombie is defeated. In this case, the Discoverer categorizes this increase as a result of defeating the zombie. For the in-domain hallucination example, it is impossible to increase the wood while defeating the zombie, though the wood may appear in the observation. Lastly, the out-domain hallucination is responsible for discovering dynamics that contain nonexistent objects in the observation or even the Crafter environment.
@@
Table 3: The dynamics underlying the outcome of defeating zombie
Compared to the dynamics from SPRING in Table 4 , we found that DiVE’s dynamics are not only more precise but also more informed. For example, regarding the dynamics of placing a stone, SPRING only found that placing a stone requires stones; however, DiVE found that placing a stone requires exactly one stone and the precise facing condition for successful placement. Furthermore, based on this information, the Evolver can derive advanced dynamics for placing a stone, such as its potential to act as a barrier between the agent and dangerous creatures.
Methods               Dynamics
SPRING Place stone requires stones
D&V Place stone requires 1 stone and faces paths, grass, sand, water, and lava
D&V&E Place stone to block zombies and skeletons, preventing them from reaching the player
Table 4: Comparing SPRING and DiVE on place stone’s dynamics
@@ 5 Related Work
Embodied Agent   Reinforcement learning (RL) has long been an effective method for developing embodied agents through trial and error Moon et al. (2024); Hafner et al. (2023); Schulman et al. (2017); Jiang et al. (2022); Hafner et al. (2020); Hessel et al. (2018); Sekar et al. (2020); Weir et al. (2022). With recent advancements in LLMs Achiam et al. (2023); Team et al. (2023); Brown et al. (2020); Touvron et al. (2023); Anil et al. (2023), their reasoning and planning capabilities have emerged as an important component for the embodied agent Wang et al. (2023a, b); Ma et al. (2023); Liang et al. (2023); Brohan et al. (2023); Nottingham et al. (2023); Silver et al. (2024); Driess et al. (2023); Shi et al. (2024); Wu et al. (2024b).
Discovering and Verifying   LLMs can uncover knowledge by inducing and deducing rules for reasoning tasks Zhu et al. (2023). Additionally, they can directly discover the underlying domain knowledge from previous trajectories and interactive experience Colas et al. (2023); Majumder et al. (2023); Fu et al. (2024a, b); Zhong et al. (2024); Zhao et al. (2024).The discovered knowledge within these methods is not hierarchically structured and mainly focuses on discovering short-horizon span knowledge. Meanwhile, due to the tendency for hallucination Zhang et al. (2023b); Dziri et al. (2022b), the knowledge discovered should be carefully verified before being used directly.
Evolving   By reflecting on past trajectories, LLMs can enhance their decision-making processShinn et al. (2024). Leveraging this capability, works such as Wang et al. (2023a); Stengel-Eskin et al. (2024); Zhang et al. (2023a); Nottingham et al. (2024) evolve new and advanced skills based on pre-defined skills. However, previous methods often assume that LLMs already possess comprehensive domain knowledge and can directly develop new insights. Additionally, they often requrie pre-defined skills in either code format or RL polices.
@@ 6 Conclusion
In this work, we present DiVE, a framework designed for long-horizon interactive planning tasks. Specifically, given a handful of demonstrations, a Discoverer can extract useful world dynamics such as an action’s preconditions and outcomes. A Verifier subsequently filters out inaccurate dynamic candidates. In an online setting, conditioned on the offline-learned knowledge, an Evolver learns situational strategies through interaction. DiVE bridges the knowledge gap between a pre-trained LLM and the target domain by autonomously learning hierarchical world dynamics and evolving contextual knowledge. Extensive experiments demonstrate its effectiveness. This work may pave the way for developing LLM-based frameworks to achieve the mutual enhancement of dynamics learning and policy learning without extensive human annotation.
@@
Experimenting with DiVE in a stochastic yet simple environment may not adequately represent all the potential real-world scenarios that an embodied agent might encounter upon deployment. Additionally, DiVE requires collecting human expert trajectories, which can be either impossible or resource-intensive in certain situations. Finally, obtaining perfect object information might not be feasible in some environments.
@@
We do not anticipate any immediate ethical or societal impact arising from our work. In this study, we aimed to bridge the knowledge gap of LLMs towards the target domain. However, despite our efforts, using LLMs as the backbone, DiVE might still hallucinate due to the inherent tendency of LLMs to hallucinate.
Appendix
Appendix A Dynamics Example from DiVE
@@ Appendix B Trajectory Example
Appendix C Precision changes over time
@@
As shown in Figure 5, the precision of the verified dynamics $\mathcal{W}$ increases during the initial step of verification. However, as the verification process progresses, further increases in precision are not observed. This suggests that the first step effectively filters out all the unreliable dynamics.
Appendix D Verbalizer
@@
Appendix E Crafter’s Tech Tree
@@
Appendix F Dynamic discovering and verification algorithm
Input: Verbalized human demonstrations $\tilde{H}$, current element $E\in TD$ for discovering, dynamics of $E$’s parents $\{\mathcal{W}_{V}\}_{V\in\text{Parents}(E)}$
Constant: Discovery steps $s$, action space $\mathcal{A}$, object set $O$, semantic trajectory transformation function $T^{E}$, dynamic distillation function $D^{E}$, semantic sampling function $S_{E}$.
Result: $\tilde{\mathcal{W}}_{E}$
$\tilde{\mathcal{W}}_{E}\leftarrow\{\}$;
$B^{E}\leftarrow T^{E}(\tilde{H})$ ;
  /* $B^{E}$ is the semantic experience bank. */
for $k\leftarrow 1$ to $s$ do
       for $e\in E$ do
             for $x\in n.attributes$ do
                   $B_{e}^{E}\leftarrow S_{E}(n,B^{E})$;
                   $y\leftarrow D^{E}(x,B_{e}^{E},\{\mathcal{W}_{V}\})$;
                   $\tilde{\mathcal{W}}_{E}\leftarrow\tilde{\mathcal{W}}_{E}\cup\{y\}$
             end for
       end for
end for
Algorithm 1 Dynamics discovering
Input: Dynamic candidate set $\tilde{\mathcal{W}}_{E}$, current level $E\in TD$ for verification, semantic experience bank $B^{E}$ concerning $E$, dynamics of $E$’s parents $\{\mathcal{W}_{V}\}_{V\in\mathbf{Parents}(E)}$
Constant: Dynamic verifying function $V^{E}$, conflict detection function $F^{E}$, semantic sampling function $S_{E}$
Result: $\mathcal{W}_{E}$
$\mathcal{W}_{E}\leftarrow\{\}$;
$\bar{\mathcal{W}}_{E}\leftarrow\{\}$;
for $w\in\tilde{\mathcal{W}}_{E}$ do
       $B^{E}_{w}\leftarrow S_{E}(w,B^{N})$;
       if $V^{E}(w,B_{w}^{E})$ is True then
             $\bar{\mathcal{W}}\leftarrow\bar{\mathcal{W}}_{E}\cup\{w\}$
       end if
end for
for $w\in\bar{\mathcal{W}}_{E}$ do
       if $F^{E}\left(w,\bar{\mathcal{W}}_{E},\{\mathcal{W}_{V}\}\right)$ is False then
             $\mathcal{W}_{E}\leftarrow\mathcal{W}_{E}\cup\{w\}$
       end if
end for
Algorithm 2 Dynamics verfication
Appendix G Prompts
G.1 Grounding
G.1.1 Subtask Selection
Given the following details:
- Subtask description: {self._transition[’subtask’]},
- Current observation: {self._transition[’state_description’]},
For deciding whether to terminate the subtask, consider:
- The previous action, provided it was executed successfully.
Output in this format: {output_format}.
G.1.2 Online Evolve
@@ G.2 Critics for Evolved Dynamics
G.2.1 Action Selection
@@
G.3 Discover Action Dynamics
G.3.1 Materials used attribute
@@
G.3.2 Immediate objects attribute
@@
G.3.3 Facing object attribute
@@
G.3.4 Inventory tool attribute
@@ G.3.5 Outcome attribute
G.3.6 Facing object change attribute
@@
G.4 Verify Action Dynamics
G.4.1 Materials used attribute
@@
G.4.2 Immediate objects attributes
@@
G.4.3 Facing object attribute
@@
G.4.4 Inventory tool attribute
@@ G.4.5 Outcome attribute
G.4.6 Facing object change attribute
@@
G.5 Discover Objects Dynamics
@@ 3. Format your response as follows: {object_relationship_format}
G.6 Verify Objects Dynamics
@@
G.7 Discover Sub-task Dynamics
@@
G.8 Verify Sub-task Dynamics
@@
G.9 Discover Sub-goal Dynamics
@@
G.10 Verify Sub-goal Dynamics

### v2

@@ 5.4 Analysis of Learned Dynamics
The correctness of the learned and verified dynamics is classified as either correct or erroneous, with errors stemming from confounders, in-domain hallucinations, or out-of-domain hallucinations. As shown in Table 5, an example of a confounder-related mistake occurs in Crafter when a simultaneous increase in health points is incorrectly attributed to the act of defeating a zombie. In this scenario, the Discoverer misclassifies the health increase as a direct result of defeating the zombie. In the case of in-domain hallucinations, the Discoverer incorrectly associates an increase in wood with defeating the zombie, even though it is not possible for wood to increase during this event, despite its presence in the observation. Lastly, out-of-domain hallucinations involve the discovery of dynamics that reference nonexistent objects in the observation or elements not present in the Crafter environment.
@@
Table 5: The dynamics underlying the outcome of defeating zombie in the Crafter environment.
Compared to the dynamics from the game manual, as shown in Table 6, we found that DiVE’s dynamics are not only more precise but also more detailed. For instance, while SPRING only identified that placing a stone requires stones, DiVE determined that it specifically requires exactly one stone and the precise facing condition needed for successful placement. Moreover, using this information, the Evolver can infer advanced dynamics for placing a stone, such as its potential to serve as a barrier between the agent and dangerous creatures.
Sources               Dynamics
Manual Place stone requires stones
Discoverer Place stone requires 1 stone and faces paths, grass, sand, water, and lava
Evolver Place stone to block zombies and skeletons, preventing them from reaching the player
Table 6: Comparing SPRING and DiVE on place stone’s dynamics in the Crafter environment.
In a customized MiniHack Lava Crossing setting, we modified the task such that it requires the agent to use the Wand of Death to freeze the lava for crossing, even though the wand’s original purpose is to zap monsters. Based on the demonstration, the Discoverer mistakenly identifies the precondition for using the wand as The player must be adjacent to a wall or obstacle that can be destroyed or altered by the wand of death, with the outcome being The wand of death was used, altering the environment by removing walls and revealing a dark area to the east, southeast, and south. This unreliable dynamic is successfully identified and filtered out by the Verifier, highlighting its crucial role in maintaining the reliability of learned dynamics.
@@ 6 Related Work
Language Models
Language models (LLMs) are trained autoregressively in a left-to-right sequence, predicting each token based on its preceding context from an internet-scale corpus. Through this training, LLMs develop a comprehensive understanding of both language and the world it represents Achiam et al. (2023); Touvron et al. (2023); Dubey et al. (2024); The (2024), enabling them to perform competently across a wide range of tasks Yao et al. (2024); Shinn et al. (2024).
Figure 4: Precision of learned dynamics before and after verification in the Crafter environment.
Embodied Agent
Building an embodied agent using LLMs is challenging because LLMs lack embodied experience Valmeekam et al. (2022); Wang et al. (2024) in downstream environments Weir et al. (2022); Shridhar et al. (2020); Côté et al. (2019). However, LLMs can still provide foundational world knowledge that serves as a prior for the agent Shi et al. (2024); Colas et al. (2023); Zhong et al. (2024); Fu et al. (2024a). A notable distinction of our work is that we do not assume LLMs have the necessary knowledge to solve tasks in specific domains.
Discover Dynamics
LLMs can discover knowledge by inducing and deducing rules for reasoning tasks Zhu et al. (2023) and by extracting underlying domain knowledge from prior trajectories and interactive experiences Colas et al. (2023); Majumder et al. (2023); Fu et al. (2024b); Zhao et al. (2024). However, the knowledge obtained through these methods is often unstructured, not suited for addressing long-horizon planning problems, and lacks verification for reliability, as it overlooks LLMs’ tendency to hallucinate Zhang et al. (2023b).
Evolve Dynamics
LLMs can refine their decision-making process by reflecting on past trajectories Shinn et al. (2024). Leveraging this ability, studies such as Wang et al. (2023); Stengel-Eskin et al. (2024); Zhang et al. (2023a); Nottingham et al. (2024) focus on evolving new and advanced skills from pre-defined ones. However, these approaches often assume that LLMs already possess comprehensive domain knowledge and can derive new dynamics based solely on their understanding.
@@ 7 Conclusion
We introduce DiVE, a framework that bridges the knowledge gap between LLMs and downstream domains. The Discoverer extracts world dynamics, while the Verifier filters unreliable candidates. In an online setting, the Evolver reasons strategies through interaction. Our experiments demonstrate that DiVE effectively bridges the knowledge gap between LLMs and downstream domains.
@@
Evaluating DiVE in embodied environments may not fully capture the diversity and complexity of real-world dynamics that an agent might encounter in practice. If the environment’s dynamics change across episodes, the offline-learned dynamics from DiVE may struggle to adapt, introducing potential biases in the agent’s understanding and hindering the development of advanced strategies. Moreover, acquiring these dynamics from human demonstrations can be difficult or impractical in certain scenarios.
@@
We do not anticipate any immediate ethical or societal impact from our work. This study aims to bridge the knowledge gap between LLMs and the target domain. However, despite our efforts, DiVE may still exhibit hallucinations due to the inherent tendency of LLMs to hallucinate.
Acknowledgements
This work is supported by the Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant (RGPIN-2021-03115).
Appendix A MiniHack
This section offers further details on the MiniHack environment, along with a comprehensive description of the prompts utilized by DiVE within this environment.
A.1 Configuration
For this environment, we employ the gpt-4o-2024-08-06 model.
A.2 Des-Files for Each MiniHack Task
In this subsection, we present the Des-Files utilized for the Lava Crossing, Wand of Death, and Quest tasks.
A.2.1 Lava Crossing
MAZE: "mylevel",
FLAGS:hardfloor
INIT_MAP: solidfill,’
GEOMETRY:center,center
MAP
|.....L.....|
|.....L.....|
|.....L.....|
|.....L.....|
|.....L.....|
ENDMAP
REGION:(0,0,12,6),lit,"ordinary"
\$left_bank = selection:fillrect (1,1,5,5)
\$right_bank = selection:fillrect (7,1,11,5)
OBJECT:(’=’,"levitation"),rndcoord(\$left_bank),blessed
STAIR:rndcoord(\$right_bank),down
BRANCH:(1,1,5,5),(0,0,0,0)
A.2.2 Wand of Death
MAZE: "mylevel",
FLAGS:hardfloor
INIT_MAP: solidfill,’
GEOMETRY:center,center
MAP
ENDMAP
REGION:(1,1,5,5),lit,"ordinary"
REGION:(6,1,26,1),lit,"ordinary"
REGION:(26,1,27,1),lit,"ordinary"
\$safe_room = selection:fillrect(1,1,5,5)
OBJECT:(’/’,"death"),(1,1),blessed
MONSTER:(’H’, "minotaur"), (26,1)
STAIR:(27,1),down
BRANCH:(1,1,5,5),(1,1,2,2)
A.2.3 Quest
MAZE: "mylevel",
FLAGS:hardfloor
INIT_MAP: solidfill,’
GEOMETRY:center,center
MAP
|.....L......|-- |-----|
|.....L........|------|.....|
|.....L.....................|
|.....L........|------|.....|
|.....L......|-- |-----|
ENDMAP
REGION:(0,0,28,6),lit,"ordinary"
\$left_bank = selection:fillrect (1,1,5,5)
\$right_bank = selection:fillrect (7,1,11,5)
\$goal_room = selection:fillrect (25,2,27,4)
OBJECT:(’/’,"cold"),(2,3),blessed
BRANCH:(1,1,3,3),(0,0,0,0)
MONSTER:random, (25,3)
STAIR:(26,3),down
A.3 Prompts for MiniHack Tasks
This section includes the prompts used to solve the MiniHack tasks.
A.3.1 Offline Dynamics Learning
The prompts for offline learning the dynamics from demonstrations.
Discover Action Dynamics
The player is playing a MiniHack game and would like you to analyze the games dynamics. Specifically, they need help identifying the general preconditions and outcomes of actions by comparing consecutive observations.
- In the previous state, the player observed: {previous_state[’state_description’]} and took the action: {action}.
- In the current state, the player observed: {state[’state_description’]}.
Definitions:
- **General precondition**: The condition that must be met for the action to succeed, regardless of the specific game state.
- **General outcome**: The effect of the action, regardless of the specific game state.
Here is the action-to-text mapping: {state[’action_prompt’]}.
Discover Object Dynamics
The player is playing a MiniHack game and would like you to analyze the games dynamics. Specifically, they need help identifying the general preconditions and outcomes of actions by comparing consecutive observations.
- In the previous state, the player observed: {previous_state[’state_description’]} and took the action: {action}.
- In the current state, the player observed: {state[’state_description’]}.
Definitions:
- **General precondition**: The condition that must be met for the action to succeed, regardless of the specific game state.
- **General outcome**: The effect of the action, regardless of the specific game state.
Here is the action-to-text mapping: {state[’action_prompt’]}.
Discover Subtask Dynamics
The player is playing a MiniHack game and would like your help in discovering the detailed steps to complete a subtask, and the preconditions and outcomes of the subtask. The player has provided the following information:
- In previous states, the player observed: {previous_state_description} and took the following actions: {previous_action}.
Definition:
- **Steps for completing the subtask**: The general and detailed sequence of actions, represented in text, that the player must follow to achieve the subtask. These steps should be general and not tied to the specific state of the environment.
- **Preconditions**: The general conditions that must be met for the subtask to be completed.
- **Outcomes**: The general effects of completing the subtask.
Here is the mapping of actions to their text representations: {state[’action_prompt’]}.
A.3.2 Online Dynamics Learning
The prompts for online evolving the dynamics from existing dynamics.
Discover Subtask Dynamics
You are playing the MiniHack game.
Your current subtask is: {subtask}.
Your current observation is: {observation}.
The primitive dynamics of the game are: {action_dynamics}.
The object dynamics of the game are: {object_dynamics}.
You are asked to evolve useful strategies that aid in completing the subtask based on the provided information.
Instructions for evolving strategies:
- The evolved dynamics should not contradict the primitive dynamics.
Now, consider:
- All the potential difficulties that may arise, based on the primitive dynamics and current observation.
- All the potential dangers that may be encountered, based on the primitive dynamics and current observation.
- The strategies should be general that do not include specific actions, directions and dis
List all the potential difficulties, dangers and the evolved strategies to resolve them.
A.3.3 Grounding
The prompts for grounding the agent.
Task Termination
You are playing the MiniHack game.
Your observation is: {description}.
Your current subtask is: {current_subtask}.
Your previous history is: {previous_history}.
The action and character mapping is: {action_prompt}.
You are asked to decide whether the current subtask should been terminated:
- If the termination conditions are met or the player is in danger or the player is in deadlock, the subtask should be terminated.
Output yes if the subtask has been completed, and no if it has not been completed.
Task Termination
You are playing the MiniHack game.
Your observation is: {description}.
Your current subtask is: {current_subtask}.
Your previous history is: {previous_history}.
The action and character mapping is: {action_prompt}.
You are asked to decide whether the current subtask should been terminated:
- If the termination conditions are met or the player is in danger or the player is in deadlock, the subtask should be terminated.
Output yes if the subtask has been completed, and no if it has not been completed.
Task Selection
You are playing the MiniHack game.
Your goal is: {task_description}.
Your observation is: {description}.
Your previously completed subtask is: {previous_subtask}.
All the subtasks are: {subtask_dynamics}.
All the environment dynamics are: {action_dynamics}.
All the object dynamics are: {object_dynamics}.
You are asked to select the next best subtask for completing the goal, based on all the provided information.
Action Selection
You are playing the MiniHack game.
Your current subtask is: {subtask}.
Your current observation is: {description}.
You also observe the feedback from the environment for the previous action: {observation}, which you need to consider for the next action.
The action and character mapping is: {action_prompt}.
Here are the strategies for the subtask: {evolved_strategy}.
Here are the primitive dynamics of the game: {action_dynamics} and {object_dynamics}.
Your past action and thought history: {history}
Now, you are asked to:
- List the potential difficulties and dangers that may arise based on the **current observation** and primitive dynamics for completing the subtask.
- Output the thoughts on the future actions how to resolve the difficulties and dangers for completing the subtask.
- Select the best action for the next step based on the thoughts.
- Do not copy the previous thoughts.
In the game, the distance is only measured as very far > far > near
Appendix B Crafter
This section provides a detailed overview of the Crafter environment and a comprehensive description of the prompts used by DiVE.
B.1 Configuration
For this setup, we utilize the gpt-4o-2024-05-13 model in the Crafter environment with seeds 1, 2, 3, 4, and 5.
B.2 An Example of Learned Dynamics
@@ B.3 An Example of Trajectory
B.4 Changes in Dynamics Precision Over Time
@@
As shown in Figure 5, the precision of the verified dynamics $\mathcal{W}$ increases significantly during the initial verification step. However, as the process continues, further improvements in precision are minimal, indicating that the first step successfully filters out most of the unreliable dynamics.
B.5 An Example of Verbalized Observation
@@
B.6 Tech Tree in the Crafter Environment
@@
B.7 Prompts for Crafter Environment
This section includes the prompts used in the Crafter environment.
B.7.1 Online Dynamics Learning
@@
B.7.2 Grounding
The prompts for grounding the agent.
Subtask Termination
Given the following details:
- Subtask description: {self._transition[’subtask’]},
- Current observation: {self._transition[’state_description’]},
For deciding whether to terminate the subtask, consider:
- The previous action, provided it was executed successfully.
Output in this format: {output_format}.
Action Selection
@@ B.7.2 Grounding
B.7.3 Offline Dynamics Learning
Learning Basic Attribute
@@ B.7.3 Offline Dynamics Learning
Learning Basic Attribute
@@
Learning Basic Attribute
@@
Learning Basic Attribute
@@
Learning Basic Attribute
@@
Verifying Basic Attribute
@@
Verifying Basic Attribute
@@
Verifying Basic Attribute
@@
Verifying Basic Attribute
@@
Verifying Basic Attribute
@@
Verifying Basic Attribute
@@ 3. Format your response as follows: {object_relationship_format}
Verifying Basic Attribute
@@
Discover Subtask Attribute
@@
Verify Subtask Attribute
@@
Discover Subgoal Attribute
@@
Verify Subgoal Attribute